Circular No. 8299

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004an AND 2004ao Further to IAUC 8297, D. Singer and W. Li report the LOSS discovery

of two apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004an Mar. 7.50 14^{h}40^{m}19.^{s}57 +16^{o}41′03′′.8 18.1 1′′.1 E, 2′′.7 S
2004ao Mar. 7.54 17 28 09.35 + 7 24 55.5 14.9 6′′.3 E, 23′′.8 S

Additional KAIT magnitudes: SN 2004an in IC 4483, Feb. 2.49 UT, [19.0; Mar. 6.54, 18.1. SN 2004ao in UGC 10862, 2003 Sept. 13.21, [19.0; 2004 Mar. 8.54, 15.0.

SUPERNOVA 2004am IN M82 S. Mattila, Stockholm Observatory; W. P. S. Meikle, Imperial Col-

lege, London; P. Groeningsson, Stockholm Observatory; R. Greimel and M. Schirmer, Isaac Newton Group of Telescopes; and J. A. Acosta-Pulido, Instituto de Astrofisica de Canarias, obtained infrared images and spectra of SN 2004am (IAUC 8297) on Mar. 6.9 UT, using LIRIS on the William Herschel Telescope (WHT). Preliminary reduction of the 0.89–1.53-μm spectrum shows discrete features superimposed on a significant continuum sloping slightly upwards to longer wavelengths. Emission lines of Paγ and Paβ are seen, having FWZI of ∼ 2800 km/s. It is concluded that this is a type-II supernova, but probably highly reddened within the starburst galaxy M82. The spectrum has some similarity to that of the type-IIP supernova 1995V (Fassia et al. 1998, MNRAS 299, 150) at ∼ 85 days. A preliminary magnitude of K_{s} = 12 was estimated for the supernova via image-subtraction techniques, using a reference K_{s}-band image obtained on 2002 Jan. 3.0 with the WHT (+ INGRID) as part of the ongoing ‘Nuclear Supernova Search’ campaign. Adopting the unfiltered magnitude (m_{u}) from IAUC 8297 yields a color of m_{u}–K ∼ +5. This indicates a high extinction for the supernova, as already suggested by W. Li (private comm.), corresponding to a visual extinction of ∼ 5 magnitudes.

COMET C/2001 Q4 (NEAT) Visual total-magnitude estimates: Feb. 19.04 UT, 7.2 (J. Shanklin,

Stanley, Falkland Islands, 10×50 binoculars); Mar. 1.77, 7.2 (D. A. J. Seargent, Cowra, N.S.W., 6×35 binoc.); 5.12, 7.2 (Shanklin, Rothera, Antarctica); 8.52, 7.5 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars).

2004 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green